OpenStudy (anonymous):

-3x(x^3+2x^2-x) and this one im suppose to multiply and simplify

Parth (parthkohli):

They should never use the word "simplify" instead of "expand".

OpenStudy (anonymous):

i totally agree with @ParthKohli "simplify" has no mathematical meaning. it is the lazy math teachers way to say "give me the answer i want"

OpenStudy (anonymous):

\[-3x(x^3+2x^2-x)=-3x\times x^3-3x\times 2x^2-3x\times (-x)\] \[=-3x^4-6x^3+3x^2\] it is the distributive law at work

Parth (parthkohli):

If you want it the hearts and spades style,\[\heartsuit ( \clubsuit + \spadesuit+ \diamondsuit ) = \heartsuit\clubsuit + \heartsuit \spadesuit + \heartsuit\diamondsuit \]